# Gaming on Coincident Peak Shaving: Equilibrium and Strategic Behavior

### v4

@@ VIII Discussion and Conclusion
  • Both agents are (lower) non-capable, i.e., $-r_{\mathrm{x}}>b_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$, and the NE is their critical point $x^{*}=-r_{\mathrm{x}},y^{*}=-r_{\mathrm{y}}$.
(1b) $X_{1}\geq X_{2},Y_{1}<Y_{2}$, i.e., agent x’s demand is higher in the CP period, while agent y’s demand is lower. Similar to (1a), when both agents are capable, the NE is their balance point. When they are not all capable agents, their optimal strategy is determined by whether they can change the CP period together and whether they can balance their own demand. Suppose agent x is non-capable, due to $X_{1}\geq X_{2}$, agent x must be lower non-capable, i.e., $b_{\mathrm{x}}<-r_{\mathrm{x}}$. Because the CP period is 1 and the agent x’s demand is higher in period 1, although agent x is lower non-capable, it is possible to change the CP period solely without balancing its own demand. Thus, we separate two scenarios to analyze.
  • Agent x can’t change the CP period solely, i.e., $-r_{\mathrm{x}}\geq b$. Agent x’s best strategy is to shift demand away from the CP period, but can’t reach the system balance point, so agent y will also shift some demand away from the CP period to reduce its cost but must compare with the shifting penalty. Thus, the NE is
    $x^{*}=-r_{\mathrm{x}},\ y^{*}=\max\{-r_{\mathrm{y}},b-x*\}.$ (35a)
    When $-r_{\mathrm{y}}>b-x^{*}$, both agents can’t change the CP period together, and the best strategy for agent y is $y^{*}=-r_{\mathrm{y}}$. The condition for agent x is $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is $b_{\mathrm{y}}<-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}$. Otherwise, $-r_{\mathrm{y}}\leq b-x^{*}$, agent y can cooperate with agent x to change the CP period, and the best strategy for agent x is $y^{*}=b-x^{*}=b+r_{\mathrm{x}}$. The condition for agent x is the same as the lower non-capable condition $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is
    $-r_{\mathrm{x}}\geq b,\ -r_{\mathrm{x}}-r_{\mathrm{y}}\leq b,$ (35b)
    $-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}\leq b_{\mathrm{y}}\leq-r_{\mathrm{x}}-b_{\mathrm{x}}.$ (35c)
  • Agent x can change the CP period solely, i.e., $-r_{\mathrm{x}}<b$. Agent x shifts demand away from the CP period and at least changes the CP period, and agent y will move demand back to the CP period to keep the original CP period. Thus, determined by which agent can shift more before reaching their critical point limit, the NE is
    $x^{*}=\max\{-r_{\mathrm{x}},b-y^{*}\},\ y^{*}=\min\{r_{\mathrm{y}},b-x^{*}\}.$ (36a)
    When $r_{\mathrm{y}}\leq b-x^{*}$, which means agent y first reaches the critical point limit, then agent x must satisfy $b-y^{*}>-r_{\mathrm{x}}$. Thus, the best strategy is $x^{*}=b-r_{\mathrm{y}},y^{*}=r_{\mathrm{y}}$. The condition for agent x is the same as the lower non-capable condition $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is
    $b-r_{\mathrm{y}}>-r_{\mathrm{x}},\ b_{\mathrm{y}}>r_{\mathrm{y}}-r_{\mathrm{x}}-b_{\mathrm{x}}.$ (36b)
    Otherwise, we have the best strategy for both agents as $x^{*}=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}}$, and the corresponding condition for agent x is still $b_{\mathrm{x}}<-r_{\mathrm{x}}$, and for agent y is
    $-r_{\mathrm{x}}-b_{\mathrm{x}}<b_{\mathrm{y}}\leq r_{\mathrm{y}}-r_{\mathrm{x}}-b_{\mathrm{x}}.$ (36c)
@@
If agent y is non-capable, due to $Y_{1}<Y_{2}$, agent y must be upper non-capable, i.e., $b_{\mathrm{y}}>r_{\mathrm{y}}$. We focus on the conditions that both agents can change the CP period together because we have shown the condition that they can’t change the CP period above. Thus, the NE is determined by which agents can shift more before reaching their critical point limit, and we have the solution structure
$y^{*}=\min\{r_{\mathrm{y}},b-x^{*}\},x^{*}=\max\{-r_{\mathrm{x}},b-y^{*}\}.$ (37a)
Following the same process with (36), we know when $-r_{\mathrm{x}}\geq b-y^{*}$, agent y must satisfy $r_{\mathrm{y}}>b-x^{*}$, and the best strategy for both agents are $x^{*}=-r_{\mathrm{x}},y^{*}=b+r_{\mathrm{x}}$. The condition for agent y is the upper non-capable condition $b_{\mathrm{y}}>r_{\mathrm{y}}$, and for agent x is
$r_{\mathrm{y}}\geq b+r_{\mathrm{x}},\ -r_{\mathrm{x}}-r_{\mathrm{y}}\leq b,$ (37b)
$-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{y}}\leq b_{\mathrm{x}}\leq-r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}},$ (37c)
where the second condition is because both agents can change the CP period together. Otherwise, the best strategy for both agents is $x^{*}=b-r_{\mathrm{y}},y^{*}=r_{\mathrm{y}}$, and the condition for agent y is still the upper non-capable condition $b_{\mathrm{y}}>r_{\mathrm{y}}$, and for agent x is
$-r_{\mathrm{x}}\leq b-r_{\mathrm{y}},\ r_{\mathrm{x}}+r_{\mathrm{y}}\geq b,$ (37d)
$-r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}.$ (37e)
(1c) $X_{1}<X_{2},Y_{1}\geq Y_{2}$, i.e., agent y’s demand is higher in the CP period, while agent x’s demand is lower, which is the same to (1b) by changing the x to y, and we omit the redundant math here.
(2) $S_{\mathrm{b},1}<S_{\mathrm{b},2}$, i.e., baseline CP period 2. Similar to the CP period 1 condition, we can still separate three cases with (2a) $X_{1}<X_{2},Y_{1}<Y_{2}$; (2b) $X_{1}\geq X_{2},Y_{1}<Y_{2}$; and (2c) $X_{1}<X_{2},Y_{1}\geq Y_{2}$. The only difference in all cases is that when the best strategy is the critical point, the critical point changes from ±$\pm r_{\mathrm{x}}$ to $\mp r_{\mathrm{x}}$. Other analysis is the same as (1), and we omit the redundant part.
To conclude, combining all the scenarios, we obtain the NE and corresponding conditions as (7) show and prove this Theorem. ∎
@@
The key is to show the rate of change of $\lVert F_{1}(x,y)\rVert^{2}$ is always negative for $F_{1}(x,y)\neq 0$ [Rosen, 1965]. We have
@@
Thus, $\lim_{k\rightarrow\infty}\lVert F_{1}\rVert=0$, so that $(x(k),y(k))\rightarrow(x^{*},y^{*})$, where $(x^{*},y^{*})$ is the equilibrium point and $F_{1}(x^{*},y^{*})=0$. Following the interior trajectory theorem from [Rosen, 1965], we know $(x^{*},y^{*})\in\mathcal{X}$, which proves this Lemma by showing system (8a) is asymptotically stable.
@@
We prove the Theorem based on the multiple Lyapunov function method [Liberzon, 2003]. Since our system dynamics is liner, the basic idea is to (i) find the functions $\mathcal{V}_{j}>0$ in each system (period) $j=1,2$, for $\dot{\mathcal{V}}_{j}\neq 0$, the function $\mathcal{V}_{j}$ is always decreased along the solution of the $j$th system in the region where this system is active, i.e., it is peak; (ii) on the switching surface $S_{1}(x,y)=S_{2}(x,y)$ the function $\mathcal{V}_{j}$’s value match.
@@
Now, let’s analyze the convergent equilibrium point. If the game is concave, only one period could be active, i.e., period 1 is active if $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$ and vice versa. Then the function value $\mathcal{V}_{j}$ decreases over the period until $\dot{\mathcal{V}}_{j}=0$, and the system reaches the stable points. According to the Theorem 5, the equilibrium point is unique and as described in (10b) or (10c).
Otherwise, both periods will be active sequentially, and $\mathcal{V}_{j}$ will decrease over time until $\mathcal{V}_{1}=\mathcal{V}_{2}$. The reason is that the dynamics from one period always push the other period active, e.g., $F_{1}$ in period 1 always decreases $x,y$, which pushes the solution past the switching surface and active period 2. Thus, as $\mathcal{V}_{1}=\mathcal{V}_{2}$ on the switching surface, both periods are finally stable on the switching surface. This means $(x(k),y(k))\rightarrow(x^{*},y^{*})$ when $k\rightarrow\infty$, where $(x^{*},y^{*})$ is the equilibrium point. According to the Theorem 5, the equilibrium point is unique and should satisfy (10a). Thus, we prove the global uniform asymptotically stable of the overall system in (11), and the equilibrium point satisfies (10).
@@
When selecting the learning rate, we gradually reduce $\tau_{\mathrm{x},h}$ by $\beta_{2}\tau_{\mathrm{x},h},\beta_{2}\in[0,1]$ until (46) satisfy. If the next step and current step lie on the same CP period, this condition ensures the objective $-f_{\mathrm{x},j}(x_{h})$ reduces by at least $\beta_{1}\tau_{\mathrm{x},h}\lVert F_{j}(x_{h},y_{h})\rVert^{2}$. This proves the concave game convergence as all the steps lie in one period, and the objective function is concave, thus, the objective function gradually reduces until $\lVert F_{j}(x_{h},y_{h})\rVert=0$.
In terms of quasiconcave and non-concave games, there are switches during the algorithm iteration. It is easy to imagine that each agent shifts demand away from the baseline CP period at the beginning, monotonically reducing their costs, and the cost function from their baseline CP period. Once they reach the balance point (switching surface), $x_{h}+y_{h}=b$, the solution starts switching between two periods. Noted that when switching happens, both agents’ individual peak period must be different, i.e., if $X_{1}+x_{h}>X_{2}-x_{h}$ for agent x, then $Y_{1}+y_{h}<Y_{2}-y_{h}$ must hold for agent y; otherwise, there will be no switching;
@@
This proves the Theorem by showing $-f_{\mathrm{x},1},-f_{\mathrm{y},2}$ reduce and $-f_{\mathrm{x},2},-f_{\mathrm{y},1}$ increase if $X_{1}+x>X_{2}-x$, following Theorem 6, we know the system will converge to either the switching surface $f_{\mathrm{x},1}+f_{\mathrm{y},1}=f_{\mathrm{x},2}+f_{\mathrm{y},2}$ or the critical points $-r_{\mathrm{x}}$ ($F_{1}(x,y)=0$) when $S_{\mathrm{b},1}\geq S_{\mathrm{b},2}$ or $r_{\mathrm{x}}$ ($F_{2}(x,y)=0$) when $S_{\mathrm{b},1}<S_{\mathrm{b},2}$.
@@
We first show the optimal solution from the centralized CP shaving model as described in (14). Under the following condition, we can easily get both agents’ solution as the critical point $r_{\mathrm{x}},r_{\mathrm{y}}$ by first-order optimality condition.
$-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b,\text{if }S_{\mathrm{b},1}\geq S_{\mathrm{b},2},$ (54a)
$r_{\mathrm{x}}+r_{\mathrm{y}}\geq b,\text{if }S_{\mathrm{b},1}<S_{\mathrm{b},2}.$ (54b)
@@
$\mathcal{L}(x,y,\lambda)$ $=\pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\lambda(b-x-y),$ (54c)
$\frac{\partial\mathcal{L}}{\partial x}$ =±$=\pm\pi+2\alpha_{\mathrm{x}}x-\lambda=0,$ (54d)
$\frac{\partial\mathcal{L}}{\partial y}$ =±$=\pm\pi+2\alpha_{\mathrm{y}}y-\lambda=0,$ (54e)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=x+y=b,$ (54f)
$x$ $=\frac{\alpha_{\mathrm{y}}}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}b,y=\frac{\alpha_{\mathrm{x}}}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}b,$ (54g)
where the sign of $\pm$ is determined by the system baseline demand $S_{\mathrm{b},1},S_{\mathrm{b},2}$. Overall, the solution for the centralized model (14) is
$x$ $=\frac{\alpha_{\mathrm{y}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},y=\frac{\alpha_{\mathrm{x}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},$ if $\text{if }-r_{\mathrm{x}}-r_{\mathrm{y}}\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}$ (55a)
$x$ $=-r_{\mathrm{x}},y=-r_{\mathrm{y}},$ if $\text{if }-r_{\mathrm{x}}-r_{\mathrm{y}}>b$ (55b)
$x$ $=r_{\mathrm{x}},y=r_{\mathrm{y}},$ if $\text{if }r_{\mathrm{x}}+r_{\mathrm{y}}<b$ (55c)
Note that the peak shaving effectiveness defined in the theorem statement is equal to directly comparing the $x^{*}+y^{*}$ with $x_{\mathrm{cen}}^{*}+y_{\mathrm{cen}}^{*}$ because the baseline demand is the same in both models. Combine with the game solution in Theorem 5, under the concave game conditions, we have $x^{*}=x_{\mathrm{cen}}^{*}=\pm r_{\mathrm{x}},y^{*}=y_{\mathrm{cen}}^{*}=\pm r_{\mathrm{y}}$, indicating the peak shaving effectiveness at equilibrium equal to 1. Under the non-concave game and quasiconcave game conditions, it is easy to see both game model and centralized model will balance system demand, i.e., $x^{*}+y^{*}=x_{\mathrm{cen}}^{*}+y_{\mathrm{cen}}^{*}=b$, meaning that the peak shaving effectiveness at equilibrium also equal to 1. This proves the theorem.
@@
Although the game solution is different under these two conditions, we can denote it as $x^{*}$, and according to the definition of PoA (18a), the PoA is
@@
We then know $P$ will increase with $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}$, and $P$ is continuous/differentiable function regarding $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and $x^{*},y^{*}$, thus $\partial P/\partial[(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}]>0$. Note that the shifting cost is $\alpha_{\mathrm{x}}x^{*2}$ for agent x and $\alpha_{\mathrm{x}}x^{*}=\partial(\alpha_{\mathrm{x}}x^{*2})/\partial x^{*}$, thus we call it marginal shifting cost. ∎

### v5

@@ VIII Discussion and Conclusion
  • Both agents are (upper) non-capable, i.e., $b_{\mathrm{x}}>r_{\mathrm{x}},b_{\mathrm{y}}>r_{\mathrm{y}}$, the game is concave and the NE of both agents are their critical point $x^{*}=r_{\mathrm{x}},y^{*}=r_{\mathrm{y}}$.
(2) $X_{1}<X_{2},Y_{1}\geq Y_{2}$, agent x has higher individual demand in the baseline CP period, while agent y has lower demand. We again consider four scenarios.
  • Both agents are capable, i.e., $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0$, similar to the first scenario in case (1), the NE remains the same as in that scenario.
  • Agent x is upper non-capable while agent y is capable, i.e., $b_{\mathrm{x}}>r_{\mathrm{x}},-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq 0$. Similar to the second scenario in case (1), agent x’s best strategy remains $x^{*}=r_{\mathrm{x}}$, but agent y’s best strategy differs due to the baseline demand conditions, given by $y^{*}=\max\{-r_{\mathrm{y}},b-x^{*}\}$. From the conditions of this scenario, we have $-r_{\mathrm{y}}-b_{\mathrm{y}}\leq 0,r_{\mathrm{x}}-b_{\mathrm{x}}<0$, leading to $-r_{\mathrm{y}}+r_{\mathrm{x}}-b_{\mathrm{y}}-b_{\mathrm{x}}=-r_{\mathrm{y}}+r_{\mathrm{x}}-b<0$. This implies $-r_{\mathrm{y}}<b-r_{\mathrm{x}}$, so the best strategy for agent y is $y^{*}=b-r_{\mathrm{x}}$. The corresponding conditions are $b_{\mathrm{x}}>r_{\mathrm{x}},-r_{\mathrm{y}}\leq b_{\mathrm{y}}$ and $-r_{\mathrm{y}}+r_{\mathrm{x}}-b_{\mathrm{x}}<b_{\mathrm{y}}$.
  • Agent x is capable while agent y is non-capable, i.e., $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$, agent y must be lower non-capable since $Y_{1}\geq Y_{2}$, and its best strategy is to shift demand up to its critical point in period 1, so $y^{*}=-r_{\mathrm{y}}$. Similar to the analysis in the second scenario in case (1), agent x’s best strategy is determined by whether it can fully balance the system demand given agent y’s strategy, i.e., $x^{*}=\max\{-r_{\mathrm{x}},b-y^{*}\}$. Since $X_{1}<X_{2}$, we must have $x\geq 0$, thus $x^{*}=b-y^{*}=b+r_{\mathrm{y}}$, and the conditions for agent x is $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$, for agent y is $b+r_{\mathrm{y}}\geq-r_{\mathrm{x}}$, equivalent to $b_{\mathrm{y}}\geq-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}$
@@
  • Both agents are non-capable, i.e., $b_{\mathrm{x}}>r_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$. Given their baseline conditions, both agents must compete, and their best strategy is determined by which agent can shift more demand before reaching their critical point. The equilibrium strategy is thus given by
    $x^{*}=\min\{r_{\mathrm{x}},b-y^{*}\},y^{*}=\max\{-r_{\mathrm{y}},b-x^{*}\}.$ (35)
    i) If $x^{*}=r_{\mathrm{x}}$ then it must hold that $r_{\mathrm{x}}<b-y^{*}$. Suppose $y^{*}=-r_{\mathrm{y}}$, then for agent y, we require $-r_{\mathrm{y}}>b-x^{*}=b-r_{\mathrm{x}}$. This results in the conditions $r_{\mathrm{x}}<b-y^{*}=b+r_{\mathrm{y}}$ for agent x. Obviously, the condition of agents x and y are conflict. Thus, the NE for both agents is given by $x^{*}=r_{\mathrm{x}},y^{*}=b-x^{*}=b-r_{\mathrm{x}}$ with the corresponding condition $r_{\mathrm{x}}<b_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$ and $-r_{\mathrm{y}}<b-r_{\mathrm{x}}$. ii) Otherwise, if $r_{\mathrm{x}}>b-y^{*}$, then $x^{*}=b-y^{*}$. Suppose $y^{*}=b-x^{*}$, then for agent y, we require $-r_{\mathrm{y}}<b-x^{*}$, which simplifies to $r_{\mathrm{x}}>b-y^{*}=x^{*}$ for agent x and $-r_{\mathrm{y}}<b-x^{*}=y^{*}$ for agent y. This implies that agent x does not reach the critical point in period 2, and agent y does not reach the critical point in period 1. Therefore, their best strategies would be the balance points, i.e., $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$. However, this contradicts the conditions of this scenario that $b_{\mathrm{x}}>r_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$. Thus, the NE for both agents are $x^{*}=b-y^{*}=b+r_{\mathrm{y}},y^{*}=-r_{\mathrm{y}}$. The correspond conditions are $r_{\mathrm{x}}<b_{\mathrm{x}},-r_{\mathrm{y}}>b_{\mathrm{y}}$ and $,-r_{\mathrm{y}}>b-r_{\mathrm{x}}$.
To conclude, combining all the scenarios, we obtain the NE and corresponding solutions for different game structures.
The game is quasiconcave when both agents are capable, i.e., $0\leq b_{\mathrm{x}}\leq r_{\mathrm{x}}$ for agent x and $-r_{\mathrm{y}}\leq b_{\mathrm{y}}\leq r_{\mathrm{y}}$ for agent y, which aligns with (6b) as described in Proposition 1. The NE in this case is $x^{*}=b_{\mathrm{x}},y^{*}=b_{\mathrm{y}}$.
The game is concave if the sum of agents’ shifting capacities is insufficient to change the CP period, which is the union of the following
$r_{\mathrm{x}}<b_{\mathrm{x}},\ r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{x}}<b_{\mathrm{y}}\leq r_{\mathrm{y}},$ (36a)
$b_{\mathrm{y}}>r_{\mathrm{y}},\ r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{y}}<b_{\mathrm{x}}\leq r_{\mathrm{x}},$ (36b)
$r_{\mathrm{x}}<b_{\mathrm{x}},r_{\mathrm{y}}<b_{\mathrm{y}}.$ (36c)
This is equivalent to $r_{\mathrm{x}}+r_{\mathrm{y}}<b$ and the same to (6a) as described in Proposition 1. The NE in this case is $x^{*}=r_{\mathrm{x}},y^{*}=r_{\mathrm{y}}$.
The game is non-concave when $0\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}$. We have three cases for the NE and corresponding conditions. i) $x^{*}=r_{\mathrm{x}},y^{*}=b-r_{\mathrm{x}}$, the conditions are the union of the following
$r_{\mathrm{x}}<b_{\mathrm{x}},\ b_{\mathrm{y}}\leq r_{\mathrm{y}},\ b_{\mathrm{y}}\leq r_{\mathrm{x}}+r_{\mathrm{y}}-b_{\mathrm{x}},$ (37a)
$r_{\mathrm{x}}<b_{\mathrm{x}},\ -r_{\mathrm{y}}\leq b_{\mathrm{y}}\ b_{\mathrm{y}}>r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}},$ (37b)
$r_{\mathrm{x}}<b_{\mathrm{x}},\ b_{\mathrm{y}}<-r_{\mathrm{y}},\ b_{\mathrm{y}}>r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}},$ (37c)
which is equivalent to $0\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}$ and $r_{\mathrm{x}}<b_{\mathrm{x}}$. The reason is that given this conditions, automatically, we have $-r_{\mathrm{y}}\leq b_{\mathrm{y}}$ and indicates $b_{\mathrm{y}}>r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}$ (37b) and (37c).
ii) $x^{*}=b-r_{\mathrm{y}},y^{*}=r_{\mathrm{y}}$, the conditions are $b<r_{\mathrm{x}}+r_{\mathrm{y}}$ and $r_{\mathrm{y}}<b_{\mathrm{y}}$. iii) $x^{*}=b+r_{\mathrm{y}},y^{*}=-r_{\mathrm{y}}$, the conditions are the union of the following
$b_{\mathrm{x}}\leq r_{\mathrm{x}},\ -r_{\mathrm{y}}>b_{\mathrm{y}},\ b_{\mathrm{y}}\geq-r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}},$ (38a)
$r_{\mathrm{x}}<b_{\mathrm{x}},\ -r_{\mathrm{y}}>b_{\mathrm{y}},\ b_{\mathrm{y}}\leq r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}},$ (38b)
which is equivalent to $0\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}$ and $-r_{\mathrm{y}}>b_{\mathrm{y}}$. Still, the reason is that given this conditions, automatically, we have $b_{\mathrm{x}}<r_{\mathrm{x}}+2r_{\mathrm{y}}$, indicating $b_{\mathrm{y}}\leq r_{\mathrm{x}}-r_{\mathrm{y}}-b_{\mathrm{x}}$ (38b)
These results are consistent with (6c) in Proposition 1 and (7c) in this theorem. This completes the proof.
@@
The key is to show the rate of change of $\lVert F_{1}(x,y)\rVert^{2}$ is always negative for $F_{1}(x,y)\neq 0$ [Rosen,, 1965]. We have
@@
Thus, $\lim_{k\rightarrow\infty}\lVert F_{1}\rVert=0$, so that $(x(k),y(k))\rightarrow(x^{*},y^{*})$, where $(x^{*},y^{*})$ is the equilibrium point and $F_{1}(x^{*},y^{*})=0$. Following the interior trajectory theorem from [Rosen,, 1965], we know $(x^{*},y^{*})\in\mathcal{X}$, which proves this Lemma by showing system (8a) is asymptotically stable.
@@
We prove the Theorem based on the multiple Lyapunov function method [Liberzon,, 2003]. Since our system dynamics is liner, the basic idea is to (i) find the functions $\mathcal{V}_{j}>0$ in each system (period) $j=1,2$, for $\dot{\mathcal{V}}_{j}\neq 0$, the function $\mathcal{V}_{j}$ is always decreased along the solution of the $j$th system in the region where this system is active, i.e., it is peak; (ii) on the switching surface $S_{1}(x,y)=S_{2}(x,y)$ the function $\mathcal{V}_{j}$’s value match.
@@
Now, let’s analyze the convergent equilibrium point. If the game is concave, only the baseline CP period 2 is active. Then the function value $\mathcal{V}_{2}$ decreases over the period until $\dot{\mathcal{V}}_{2}=0$, and the system reaches the stable points. According to the Theorem 1, the stable point is the unique equilibrium point as described in (10b).
Otherwise, both periods will be active sequentially, and $\mathcal{V}_{j}$ will decrease over time until $\mathcal{V}_{1}=\mathcal{V}_{2}$. The reason is that the dynamics from one period always push the other period active, e.g., $F_{1}$ in period 1 always decreases $x,y$, which pushes the solution past the switching surface and active period 2. Thus, as $\mathcal{V}_{1}=\mathcal{V}_{2}$ on the switching surface, both periods are finally stable on the switching surface. This means $(x(k),y(k))\rightarrow(x^{*},y^{*})$ when $k\rightarrow\infty$, where $(x^{*},y^{*})$ is the stable point. According to the Theorem 1, the stable point should be the unique equilibrium point and satisfy (10a). Thus, we prove the global uniform asymptotically stable of the overall system in (11), and the equilibrium point satisfies (10).
@@
When selecting the learning rate, we gradually reduce $\tau_{\mathrm{x},h}$ by $\beta_{2}\tau_{\mathrm{x},h},\beta_{2}\in[0,1]$ until (47) satisfy. If the next step and current step lie on the same CP period, this condition ensures the objective $-f_{\mathrm{x},j}(x_{h})$ reduces by at least $\beta_{1}\tau_{\mathrm{x},h}\lVert F_{j}(x_{h},y_{h})\rVert^{2}$. This proves the concave game convergence as all the steps lie in the baseline CP period 2, and the objective function is concave, thus, the objective function gradually reduces until $\lVert F_{2}(x_{h},y_{h})\rVert=0$.
In terms of quasiconcave and non-concave games, there are switches during the algorithm iteration. It is easy to imagine that each agent shifts demand away from the baseline CP period at the beginning, monotonically reducing their costs following the cost function from their baseline CP period. Once they reach the balance point (switching surface), $x_{h}+y_{h}=b$, the solution starts switching between two periods. Noted that when switching happens, both agents’ individual peak period must be different, i.e., if $X_{1}+x_{h}>X_{2}-x_{h}$ for agent x, then $Y_{1}+y_{h}<Y_{2}-y_{h}$ must hold for agent y; otherwise, there will be no switching;
@@
This proves the Theorem by showing we are able to choose learning rate $\tau_{\mathrm{x}},\tau_{\mathrm{y}}$ that satisfy $-f_{\mathrm{x},1},-f_{\mathrm{y},2}$ reduce and $-f_{\mathrm{x},2},-f_{\mathrm{y},1}$ increase when $X_{1}+x>X_{2}-x$. Reversely, when $X_{1}+x\leq X_{2}-x$, we can also choose learning rate such that $-f_{\mathrm{x},1},-f_{\mathrm{y},2}$ increase and $-f_{\mathrm{x},2},-f_{\mathrm{y},1}$ reduce.
@@
We first show the optimal solution from the centralized CP shaving model as described in (14). Under the condition of $r_{\mathrm{x}}+r_{\mathrm{y}}\geq b$, we can easily get both agents’ solutions are their critical point, i.e., $x_{\mathrm{cen}}^{*}=r_{\mathrm{x}},y_{\mathrm{cen}}^{*}=r_{\mathrm{y}}$ by first-order optimality condition.
@@
$\mathcal{L}(x,y,\lambda)$ $=\pi\max\{S_{1}(x,y),S_{2}(x,y)\}+\alpha_{\mathrm{x}}x^{2}+\alpha_{\mathrm{y}}y^{2}+\lambda(b-x-y),$ (55a)
$\frac{\partial\mathcal{L}}{\partial x}$ =±$=\pm\pi+2\alpha_{\mathrm{x}}x-\lambda=0,$ (55b)
$\frac{\partial\mathcal{L}}{\partial y}$ =±$=\pm\pi+2\alpha_{\mathrm{y}}y-\lambda=0,$ (55c)
$\frac{\partial\mathcal{L}}{\partial\lambda}$ $=x+y=b,$ (55d)
$x$ $=\frac{\alpha_{\mathrm{y}}}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}b,y=\frac{\alpha_{\mathrm{x}}}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}}b,$ (55e)
where the sign of $\pm$ is determined by the CP period $S_{1},S_{2}$. Overall, the solution for the centralized model (14) is
$x_{\mathrm{cen}}^{*}$ $=\frac{\alpha_{\mathrm{y}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},y_{\mathrm{cen}}^{*}=\frac{\alpha_{\mathrm{x}}b}{\alpha_{\mathrm{x}}+\alpha_{\mathrm{y}}},$ if $\text{if }0\leq b\leq r_{\mathrm{x}}+r_{\mathrm{y}}$ (56a)
$x_{\mathrm{cen}}^{*}$ $=r_{\mathrm{x}},y_{\mathrm{cen}}^{*}=r_{\mathrm{y}},$ if $\text{if }r_{\mathrm{x}}+r_{\mathrm{y}}<b$ (56b)
Note that the peak shaving effectiveness defined in the theorem statement is equal to directly comparing the $x^{*}+y^{*}$ with $x_{\mathrm{cen}}^{*}+y_{\mathrm{cen}}^{*}$ because the baseline demand is the same in both models. Combine with the game solution in Theorem 1, under the concave game conditions, we have $x^{*}=x_{\mathrm{cen}}^{*}=r_{\mathrm{x}},y^{*}=y_{\mathrm{cen}}^{*}=r_{\mathrm{y}}$, indicating the peak shaving effectiveness at equilibrium equal to 1. Under the non-concave game and quasiconcave game conditions, it is easy to see both game model and centralized model will balance system demand, i.e., $x^{*}+y^{*}=x_{\mathrm{cen}}^{*}+y_{\mathrm{cen}}^{*}=b$, meaning that the peak shaving effectiveness at equilibrium also equal to 1. This proves the theorem.
@@
Although the game solution is different under these two conditions, we can denote it as $x^{*}$, and according to the definition of efficiency loss (18a), we have
@@
We then know $P$ will increase with $(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}$, and $P$ is continuous/differentiable function regarding $\alpha_{\mathrm{x}},\alpha_{\mathrm{y}}$ and $x^{*},y^{*}$, thus $\partial P/\partial[(\alpha_{\mathrm{x}}x^{*}-\alpha_{\mathrm{y}}y^{*})^{2}]>0$. Note that the shifting cost is $\alpha_{\mathrm{x}}x^{*2}$ for agent x and $\alpha_{\mathrm{x}}x^{*}=\partial(\alpha_{\mathrm{x}}x^{*2})/\partial x^{*}$ is the marginal shifting cost. ∎